the Area of a right triangle is 24x^3+8x^2-12x-4. which of the following could be the base and the height ?
a)(4x^2+1)(3x-2)
b)(4x^2-2)(3x+2)
C) (4x^2+2)(3x-1)
D) (4x^2-2) (3x-1)

Answer by richwmiller(17219)

None of them.
Use FOIL to multiply
1/2 b*h=24x^3+8x^2-12x-4
a) (4x^2+1)(3x-2)=12x^3-8x^2+3x-2
b) (4x^2-2)(3x+2)=12x^3+8x^2-6x-4
C) (4x^2+2)(3x-1)=12x^3-4x^2+6x-2
D) (4x^2-2)(3x-1)=12x^3-4x^2-6x+2
We want half of those amounts not double.
Only b has the right signs.
Which one works out to 24x^3+8x^2-12x-4
Remember you take 1/2 of b*h